OpenStudy (anonymous):

MEDAL AND FAN WILL BE REWARDED!!!!!!!!!!!!! One leg of a right triangle is 18 in. long, and the hypotenuse is 82 in. long. Find the length of the other side of the right triangle. If necessary, round to the nearest hundredth.

OpenStudy (anonymous):

is it multiple answers?

OpenStudy (anonymous):

Nope :(

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a^2 + b^2 = c^2 hence 18^2 + b^2 = 82^2 b = 78.9746795

OpenStudy (anonymous):

wait 18in is adacent or opposite?

OpenStudy (anonymous):

as you take root, it should be = +- 78.9746795 so it actually has 2 answers.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Thank you @Mathbreaker

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] \[18^2+b^2=82^2\] b = 80

OpenStudy (anonymous):

I don't know who to give a medal to lol

OpenStudy (anonymous):

the formula i used is Pythagorean theorem actually. You might want to look into that :) And welcome :)

OpenStudy (anonymous):

give it ot mathbreaker @uglythang101

OpenStudy (anonymous):

I never expected this many people lol

OpenStudy (arabpride):

ill help ya out ~ ill give a medal to who else deserved it as well *^.^*

OpenStudy (anonymous):

@uglythang101 u changed ur pic lol

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

its a nice 1 though

OpenStudy (anonymous):

thx, I have more but will close this post then post another question!

OpenStudy (anonymous):

okay